OpenStudy (anonymous):

HEELP! Find each of the following functions. f(x)=2-3x, g(x)=cosx fog gof fof gog and state the domain of the function. please explain to me step by step, not just the answer

OpenStudy (anonymous):

f(g(x)) = 2-3(cos(x)) g(f(x)) = cos(2-3x) f(f(x)) = 2-3(2-3x) = 2-6+9x = 9x-4

OpenStudy (anonymous):

thanks, and how would it be for gog? do you have any idea of how to state the domain of these functions?

OpenStudy (anonymous):

here, i will explain the concept of f(g(x)). the idea is that you substitute g(x) in for every x in f(x). therefore, g(g(x)) means you place g(x) back into wherever there is a x in g(x). so try it and post your work. I'll check it. The domain is just what x values can you put in and get an answer that is not undefined. So for f(g(x)) = 9x-4. the domain is whatever x values you can put in and get an answer that is not undefined. Therefore, that would be \[-\infty \le x \le +infinity\]